OpenStudy (anonymous):

When you simplify each side of the equation 2(3n – 5) + 4 = 15 –(9n – 4), you will get the equivalent equation 6n – 6 = -9n + 11.

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

when u simplify each side does it come out to be the one it shows

OpenStudy (anonymous):

... 2(3n – 5) + 4 = 15 –(9n – 4) 6n-10+4=15-9n+4 6n-6=-9n+19 so the answer is..? True or False?

OpenStudy (anonymous):

do you know why 19 and not 11 ?